OpenStudy (anonymous):

I will give medals to you if you get the answer correct.. Solve the system of equations. y = 3x + 5 y = x2 + 8x + 9 A.) (-1, 2) and (-4, -7) B.) (-1, -7) and (-4, 2) C.) (1, 8) and (4, -7) D.) (-1, -8) and (-4, -17)

OpenStudy (anonymous):

The problem to solve is: \[y = 3x + 5\]\[y = x2 + 8x + 9 \]The goal here is to solve those two equations for the variables y and x. The solution to these equations would be: \[y = \frac{ 23 }{ 7 }\]\[x =- \frac{ 4 }{ 7 }\]

OpenStudy (anonymous):

Are you sure the equation is y = x2 + 8x + 9? Or y = 2x + 8x + 9?